How many molecules of iron sulphate are in 250ml of the solution?
How many ions fe3+ does the 50ml solution contain?
1
Expert's answer
2019-12-05T08:19:52-0500

Solution.

1.

"\frac{m}{M} = \frac{N}{Na}"

"C = \frac{m}{M \times V} = \frac{N}{Na \times V}"

"N = C \times Na \times V"

"N = 1 \ M \times 6.02 \times 10^{23} \times 0.25 = 1.51 \times 10^{23}"

2.

"Fe2(SO4)3 = 2Fe^{3+} + 3SO4^{2-}"

"C(Fe^{3+}) = C(Fe2(SO4)3) \times 2"

"C(Fe^{3+}) = 1 \ M \times 2 = 2 \ M"

Answer:

1.

"N = 1.51 \times 10^{23}"

2.

"C(Fe^{3+}) = 2 \ M"